### v1

@@ 3.5 Estimate the approximation rate
(29) $\begin{split}E_{\text{rand}}^{\,\text{sub}}:=&\,\max_{x\in[0,1]}\left|f_{\text{sub}}^{\text{NN}}(x)-f^{*}(x)\right|\\ \leq&\,\max_{x\in[0,1]}\left|f_{\text{sub}}^{\text{NN}}(x)-f_{\text{equi}}^{\text{NN}}(x)\right|+\max_{x\in[0,1]}\left|f_{\text{equi}}^{\text{NN}}(x)-f^{*}(x)\right|<\varepsilon/2.\end{split}$
c) Estimate the probability $P_{\text{sub}}$ related to the approximation power of $f_{\text{sub}}^{\text{NN}}$. Here we estimate the probability of finding such a subnetwork $f_{\text{sub}}^{\text{NN}}$ with the required approximation error $E_{\text{rand}}^{\,\text{sub}}$ in Eq. (29). We start with the complement event $\mathcal{A}_{k}^{c}$: for a given location $\hat{b}_{k}\in B_{\text{equi}}^{(\hat{n})}$, there is no close enough locations in $B_{\text{rand}}^{(n)}$, i.e., falls in the interval $[\hat{b}_{k}-\Delta r,\hat{b}_{k}+\Delta r]$. Concretely, the probability has $\mathbb{P}[\mathcal{A}_{k}^{c}]=(1-2\Delta r)^{n}$ since the interval length is $2\Delta r$. Hence, by choosing $\Delta r$ small enough such that these intervals have no overlap, i.e., $\Delta r<\min\{\delta_{s},1/2\hat{n}\}$, we apply the inclusion-exclusion principle [15] to write the probability of finding all locations $B_{\text{equi}}^{(\hat{n})}$ in $B_{\text{rand}}^{(n)}$ as
$\begin{split}&\mathbb{P}\left[\,\bigcap_{k=1}^{\hat{n}}\mathcal{A}_{k}\,\right]=1-\mathbb{P}\left[\,\bigcup_{k=1}^{\hat{n}}\mathcal{A}_{k}^{c}\,\right]=1-\sum_{k=1}^{\hat{n}}(-1)^{k+1}\binom{\hat{n}}{k}\left(1-2k\Delta r\right)^{n}=:1-P^{\prime},\end{split}$
where $\binom{\hat{n}}{k}$ is the binomial coefficient, and $P^{\prime}$ is the complement probability. This probability also holds for finding $W_{\text{equi}}^{(2\hat{n})}$ in pairwise $W_{\text{rand}}^{(2n)}$. Consequently, we can find a subnetwork $f_{\text{sub}}^{\text{NN}}$ within $f_{\text{rand}}^{\text{NN}}$ that is close enough to $f_{\text{equi}}^{\text{NN}}$, such that the approximation error $E_{\text{rand}}^{\,\text{sub}}$ in Eq. (29) is achieved with the probability of
(30) $P_{\text{sub}}=\mathbb{P}\big{[}E_{\text{rand}}^{\,\text{sub}}<\varepsilon/2\big{]}=\left[1-P^{\prime}\right]^{2}.$
For given $\hat{n}$ and $\Delta r$, we have $P^{\prime}\to 0$ as $n\to\infty$. Therefore, we choose a sufficiently large $n$ to ensure that the probability $P_{\text{sub}}\geq\sqrt{1-\delta}$.
d) Annihilate the remaining part in the randomly initialized $f_{\text{rand}}^{\text{NN}}$. We follow the same discussion as in the equidistant case to eliminate the unused parameters in $f_{\text{rand}}^{\text{NN}}$. By applying linear reorganization and Lemma 3.2, we rewrite the remaining part as a linear function $\mathcal{S}_{\text{un}}^{r}(x)=\beta_{r}x+\eta_{r}$, where $x\in[0,1]$ and $0\leq\beta_{r}\leq\Delta p$. The upper bound $\Delta p$ is the largest gap between the adjacent coefficients $\{p_{i}\}_{i\in I_{\text{un}}}$. As the network width $n$ increases, the randomly initialized $\{p_{i}\}_{i\in I_{\text{un}}}$ become denser, leading to $\mathbb{E}[\Delta p]\to 0$. Therefore, we can choose a sufficiently large $n$ to ensure $\Delta p\leq\varepsilon/2$ with high probability, i.e., $P_{\text{un}}\geq\sqrt{1-\delta}$, where $P_{\text{un}}$ denotes the related possibility.
@@
(31) $E_{\text{rand}}^{\,\text{un}}=\max_{x\in[0,1]}\left|\mathcal{S}_{\text{un}}^{r}(x)+C_{r}\right|\leq\Delta p\leq\varepsilon/2,\quad\text{ with probability }P_{\text{un}}>\sqrt{1-\delta}.$
@@
(32) $\left|f_{\text{rand}}^{\text{NN}}(x)-f^{*}(x)\right|\leq E_{\text{rand}}^{\,\text{sub}}+E_{\text{rand}}^{\,\text{un}}<\varepsilon,\quad\forall\,x\in[0,1].$
@@ 4.1 The algorithmic implementation of permutation training
In the implementation of permutation training, guidance is crucial in finding the weights’ ideal order relationship. The lookahead permutation (LaPerm) algorithm [34] introduces an $k$-times Adam-based free updating [25], where the learned relationship can then serve as a reference for permutation. To ensure the performance, the weights are permuted after every $k$ epoch. The impact of $k$’s value on convergence behavior is evaluated to be negligible (see Appendix C). Apart from the fixed permutation period $k$ [34], it’s also possible to adjust $k$ to learn sufficient information for the next permutation. Refer to Appendix B for more details about the LaPerm algorithm.
@@ 4.2 Experimental settings
To validate our theoretical results, we conduct experiments on regression problems. The settings are deliberately chosen to be consistent with our proof construction. We consider a three-layer network in Eq. (2), where the first hidden layer’s parameters are fixed to form the basis functions $\{\phi_{i}^{\pm}\}_{i=1}^{n}$ in Eq. (1). The weights $\theta^{(2n)}$ of the second hidden layer are trained by permutation, while the scaling factors $\alpha,\gamma$ in the output layer are freely trained to reduce the required network width. All the experiments below are repeated 10 times with different random seeds, and the error bars mark the range of the maximum and minimum values. Refer to Appendix A for the detailed experimental environment and setting.
@@ 4.4 Approximating the multi-dimensional continuous functions
As a natural extension, we consider a two-dimensional functions $z=-\sin\pi xy$, where $(x,y)\in[-1,1]^{2}$, starting with the construction of basis functions like $\phi_{i}^{\pm}(x)$ in Eq. (1). Recall that in the one-dimensional case, the two subsets of basis $\phi_{i}^{\pm}(x)$ correspond the two opposite directions along the $x$-axis. Therefore, at least four directions are required to represent a function defined on the $xy$-plane, of which two are parallel to the $x$-axis as $\phi_{i}^{\pm}(x,\cdot)=\text{ReLU}(\pm(x-b_{i}))$ and $\phi_{j}^{\pm}(\cdot,y)=\text{ReLU}(\pm(y-b_{j}))$ for $y$-axis, respectively. Furthermore, inspired by the lattice Boltzmann method in fluid mechanics [9], we introduce another four directions as $\psi_{k}^{\pm\pm}(x,y)=\text{ReLU}(\pm x\pm y-b_{k})$. So the whole basis functions are divided into eight subsets, each corresponding to a different direction (see Fig. 3(b)). Also, the range of biases is essential since the distribution of $\psi_{k}^{\pm\pm}(x,y)$ must be at least $\sqrt{2}$-times wider to cover the entire domain. Here we set the biases to range in varying directions with a uniform scaling factor, providing flexibility in dealing with the unwanted coefficients.
@@
Accordingly, we utilize a $2$-$8n$-$1$-$1$ network architecture and follow the same setting as before (refer to Appendix A). The results depicted in Fig. 3(a) also show good approximation power. However, the $1/2$ convergence rate in previous cases cannot be attained here. We hypothesize that this is due to our preliminary eight-direction setting of the basis functions. This degeneration indicates the challenge of extending our theoretical results to higher dimensions. Further research will address this difficulty by considering more appropriate high-dimensional basis function setups. One possible approach relies on the basis construction utilized in the finite element methods [3]. However, adopting such a method to meet the permutation training scenario is non-trivial and requires further investigation.
@@ 4.5 The influence of various initialization strategies
Figure 5: The performance of different initialization strategies in approximating $y=-\sin(2\pi x)$ in $[-1,1]$. The pairwise initialization $W^{(2n)}=(\pm p_{i})_{i=1}^{n},\,p_{i}\sim\mathcal{U}[-1,1]$ is noted as $W^{(2n)}\sim\mathcal{U}^{\pm}[0,1]^{n}$. The error bars are omitted for conciseness. The inset panel presents the target function as lines and an example of the approximation result as dots.
Next, we consider some commonly used initializations, such as Xavier’s uniform initialization $U_{X}$ [18], and He’s normal initialization $N_{H}$ [19]. However, the implementation of $U_{X}$ on Random 5 and $N_{H}$ on Random 6 fails to result in convergence. This abnormal poor performance may be attributed to the mismatch of the scale in $B^{(n)}$ and the approximation domain $[0,1]$. To verify our hypothesis, we hold $B^{(n)}\sim\mathcal{U}[-1,1]^{n}$, and only apply $U_{X}$ and $N_{H}$ on the coefficients $W^{(2n)}$ in Random 7 and 8. Consequently, both cases successfully regain the approximation ability. These surprising behaviors, especially the unexpected deterioration of the default choices, emphasizes the limited understanding of systematical characterizing the initialization suitable for permutation training scenarios.
@@ 4.6 Observation of the permutation-active patterns
This section aims to explore the theoretical potential of permutation training in describing network learning behavior. Based on the significant correlation between permutation and learning behavior, as evidenced by previous results [34] and our investigation, we hypothesize that the permutation-active components of the weights may play a crucial role in the training process. Therefore, by identifying and tracing the permutation-active part of weights, a novel description tool can be achieved, which also facilitates visualization and statistical analysis of the learning behavior.
@@
As a preliminary attempt, we illustrate the permutation behavior of the coefficients $\theta^{(2n)}$ in Fig. 2(a). The components that participated in the permutation are visually highlighted in dark green in Fig. 6(a). The behaviors plotted in Fig. 6(b)-(c) clearly show that the frequency of order relationship exchange evolves synchronously with the learning process, agreeing with the previous observation [34].
Specifically, the distribution of active components shows significant patterns, which are classified into four stages (marked by red dash lines in Fig. 6). The loss declines sharply in the initial stage, while only the components with medium value are permuted. Once loss reaches a plateau in the second stage, more components are involved in permutation, evidencing the role of permutation in propelling the training. As loss starts to decline again, the permutation frequency correspondingly diminishes. Interestingly, the slower loss decrease gives rise to a ribbon-like pattern, akin to the reported localized permutations (plotted in Fig. 14 of the appendix in [34]), and possibly due to slow updates failing to trigger a permutation. This observation may support the existence of inherent low-dimensional structures within the permutation training dynamics, potentially linked to mathematical depiction of permutation groups, such as cycle decomposition [6] and Fourier bases for permutation [22]. Finally, the permutation’s saturation aligns with the stationary state of loss convergence. We believe these inspiring phenomena deserve further exploration, as they hold the potential to provide novel insights regarding the learning behavior of networks.
@@ 5.1 Generalizing to other scenarios
Although we mainly focus on basic settings, the proof idea exhibits generalizability. Extending to networks equipped with leaky-ReLU can be straightforward (refer to Appendix D for numerical evidence). Our proof also enables implementations within other designs, such as networks with deeper architectures or sparse initializations (see Appendix E for detailed discussion).
However, extending our theoretical results to the high-dimensional scenario still faces challenges. One of the primary obstacles is constructing multi-dimensional basis functions that are suitable for the permutation training scenario. A reasonable approach relies on the construction in the finite element methods [3]. We plan to address this problem in future work. As for the numerical evidence of high-dimensional scenarios, the classification problem using VGG-based networks on the CIFAR-10 dataset has been examined [34], while our experiments in regression task have shown approximation behavior for two and three-dimensional inputs (see Section 4.4).
@@ 5.2 Permutation training as a theoretical tool
Our observation in Sec. 4.6 indicates the theoretical potential of permutation training, as it corresponds well with the training process and has systematical mathematical descriptions. Specifically, the patterns observed in Fig. 6 can intuitively lead to certain weight categorization strategies, potentially benefit consolidating the crucial weights for previous tasks [31], or pruning to find the ideal subnetwork in the lottery ticket hypothesis [16]. Additionally, the existing permutation training algorithm shares the same form as weight projection methods in continual learning [46], as it can be viewed as applying an order-preserving projection from the free training results to the initial weight value.
@@ 5.3 The algorithmic implementation
This work is expected to facilitate the applications of permutation training. However, some practical issues still exist and deserve further investigation. As a preliminary attempt, the existing permutation training algorithm LaPerm guides the permutation by inner loops, thus incurring undesirable external computation costs. However, employing more advanced and efficient search approaches, such as the learn-to-rank formalism [7], or permutation search algorithms in the study of LMC [24, 2], the benefits of permutation training will be actualized in practice. Importantly, our proof does not rely on any specific algorithmic implementations. Additionally, the incompatible initialization issue plotted in Fig. 2(b) emphasizes the need for developing more effective initializations as well as investigating the criterion of UAP-compatible initializations.

### v2

@@ 3.5 Estimate the approximation rate
$\begin{split}E_{\text{rand}}^{\,\text{sub}}:=&\,\max_{x\in[0,1]}\left|f_{\text{sub}}^{\text{NN}}(x)-f^{*}(x)\right|\\ \leq&\,\max_{x\in[0,1]}\left|f_{\text{sub}}^{\text{NN}}(x)-f_{\text{equi}}^{\text{NN}}(x)\right|+\max_{x\in[0,1]}\left|f_{\text{equi}}^{\text{NN}}(x)-f^{*}(x)\right|<\varepsilon/2.\end{split}$ (29)
c) Estimate the probability $P_{\text{sub}}$ related to the approximation power of $f_{\text{sub}}^{\text{NN}}$. Here we estimate the probability of finding such a subnetwork $f_{\text{sub}}^{\text{NN}}$ with the required approximation error $E_{\text{rand}}^{\,\text{sub}}$ in Eq. (29). We start with the complement event $\mathcal{A}_{k}^{c}$: for a given location $\hat{b}_{k}\in B_{\text{equi}}^{(\tilde{n})}$, there is no close enough locations in $B_{\text{rand}}^{(n)}$, i.e., falls in the interval $[\hat{b}_{k}-\Delta r,\hat{b}_{k}+\Delta r]$. Concretely, the probability has $\mathbb{P}[\mathcal{A}_{k}^{c}]=(1-2\Delta r)^{n}$ since the interval length is $2\Delta r$. Hence, by choosing $\Delta r$ small enough such that these intervals have no overlap, i.e., $\Delta r<\min\{\delta_{s},1/2\tilde{n}\}$, we apply the inclusion-exclusion principle (Feller, 1968) to write the probability of finding all locations $B_{\text{equi}}^{(\tilde{n})}$ in $B_{\text{rand}}^{(n)}$ as
$\begin{split}&\mathbb{P}\left[\,\bigcap_{k=1}^{\tilde{n}}\mathcal{A}_{k}\,\right]=1-\mathbb{P}\left[\,\bigcup_{k=1}^{\tilde{n}}\mathcal{A}_{k}^{c}\,\right]=1-\sum_{k=1}^{\tilde{n}}(-1)^{k+1}\binom{\tilde{n}}{k}\left(1-2k\Delta r\right)^{n}=:1-P^{\prime}.\end{split}$
where $\binom{\tilde{n}}{k}$ is the binomial coefficient, and $P^{\prime}$ is the complement probability. This probability also holds for finding $W_{\text{equi}}^{(2\tilde{n})}$ in pairwise $W_{\text{rand}}^{(2n)}$. Consequently, we can find a subnetwork $f_{\text{sub}}^{\text{NN}}$ within $f_{\text{rand}}^{\text{NN}}$ that is close enough to $f_{\text{equi}}^{\text{NN}}$, such that the approximation error $E_{\text{rand}}^{\,\text{sub}}$ in Eq. (29) is achieved with the probability of
$P_{\text{sub}}=\mathbb{P}\big{[}E_{\text{rand}}^{\,\text{sub}}<\varepsilon/2\big{]}=\left[1-P^{\prime}\right]^{2}.$ (30)
For given $\tilde{n}$ and $\Delta r$, we have $P^{\prime}\to 0$ as $n\to\infty$. Therefore, we choose a sufficiently large $n$ to ensure that the probability $P_{\text{sub}}\geq\sqrt{1-\delta}$.
[…]
@@
$E_{\text{rand}}^{\,\text{un}}=\max_{x\in[0,1]}\left|\mathcal{S}_{\text{un}}^{r}(x)+C_{r}\right|\leq\Delta p\leq\varepsilon/2,\quad\text{ with probability }P_{\text{un}}>\sqrt{1-\delta}.$ (31)
@@
$\left|f_{\text{rand}}^{\text{NN}}(x)-f^{*}(x)\right|\leq E_{\text{rand}}^{\,\text{sub}}+E_{\text{rand}}^{\,\text{un}}<\varepsilon,\quad\forall\,x\in[0,1].$ (32)
@@ 4.1 The algorithmic implementation of permutation training
In the implementation of permutation training, guidance is crucial in finding the weights’ ideal order relationship. Qiu and Suda (2020) proposed lookahead permutation (LaPerm) algorithm by introducing an $k$-times Adam-based free updating, where the learned relationship can then serve as a reference for permutation. To ensure the performance, the weights are permuted after every $k$ epoch. The impact of $k$’s value on convergence behavior is evaluated to be negligible (see C). Apart from the fixed permutation period $k$, it’s also possible to adjust $k$ to learn sufficient information for the next permutation. Refer to B for more details about the LaPerm algorithm.
@@ 4.2 Experimental settings
To validate our theoretical results, we conduct experiments on regression problems. The settings are deliberately chosen to be consistent with our proof construction. We consider a three-layer network in Eq. (2), where the first hidden layer’s parameters are fixed to form the basis functions $\{\phi_{i}^{\pm}\}_{i=1}^{n}$ in Eq. (1). The weights $\theta^{(2n)}$ of the second hidden layer are trained by permutation, while the scaling factors $\alpha,\gamma$ in the output layer are freely trained to reduce the required network width. All the experiments below are repeated 10 times with different random seeds, and the error bars mark the range of the maximum and minimum values. Refer to A for the detailed experimental environment and setting.
@@ 4.4 Approximating the multi-dimensional continuous functions
As a natural extension, we consider a two-dimensional functions $z=-\sin\pi xy$, where $(x,y)\in[-1,1]^{2}$, starting with the construction of basis functions like $\phi_{i}^{\pm}(x)$ in Eq. (1). Recall that in the one-dimensional case, the two subsets of basis $\phi_{i}^{\pm}(x)$ correspond the two opposite directions along the $x$-axis. Therefore, at least four directions are required to represent a function defined on the $xy$-plane, of which two are parallel to the $x$-axis as $\phi_{i}^{\pm}(x,\cdot)=\text{ReLU}(\pm(x-b_{i}))$ and $\phi_{j}^{\pm}(\cdot,y)=\text{ReLU}(\pm(y-b_{j}))$ for $y$-axis, respectively. Furthermore, inspired by the lattice Boltzmann method in fluid mechanics (Chen and Doolen, 1998), we introduce another four directions as $\psi_{k}^{\pm\pm}(x,y)=\text{ReLU}(\pm x\pm y-b_{k})$. So the whole basis functions are divided into eight subsets, each corresponding to a different direction (see Fig. 3(b)). Also, the range of biases is essential since the distribution of $\psi_{k}^{\pm\pm}(x,y)$ must be at least $\sqrt{2}$-times wider to cover the entire domain. Here we set the biases to range in varying directions with a uniform scaling factor, providing flexibility in dealing with the unwanted coefficients.
@@
Accordingly, we utilize a $2$-$8n$-$1$-$1$ network architecture and follow the same setting as before (refer to A). The results depicted in Fig. 3(a) also show good approximation power. However, the $1/2$ convergence rate in previous cases cannot be attained here. We hypothesize that this is due to our preliminary eight-direction setting of the basis functions. This degeneration indicates the challenge of extending our theoretical results to higher dimensions. Further research will address this difficulty by considering more appropriate high-dimensional basis function setups. One possible approach relies on the basis construction utilized in the finite element methods (Brenner and Scott, 2008). However, adopting such a method to meet the permutation training scenario is non-trivial and requires further investigation.
@@ 4.5 The influence of various initialization strategies
Figure 5: The performance of different initialization strategies in approximating $y=-\sin(2\pi x)$ in $[-1,1]$. The pairwise initialization $W^{(2n)}=(\pm p_{i})_{i=1}^{n},\,p_{i}\sim\mathcal{U}[-1,1]$ is denoted as $W^{(2n)}\sim\mathcal{U}^{\pm}[0,1]^{n}$. The error bars are omitted for conciseness. The inset panel presents the target function as lines and an example of the approximation result as dots.
Next, we consider some commonly used initializations, such as Xavier’s uniform initialization $U_{X}$ (Glorot and Bengio, 2010), and He’s normal initialization $N_{H}$ (He et al., 2015). However, the implementation of $U_{X}$ on Random 5 and $N_{H}$ on Random 6 fails to result in convergence. This abnormal poor performance may be attributed to the mismatch of the scale in $B^{(n)}$ and the approximation domain $[0,1]$. To verify our hypothesis, we hold $B^{(n)}\sim\mathcal{U}[-1,1]^{n}$, and only apply $U_{X}$ and $N_{H}$ on the coefficients $W^{(2n)}$ in Random 7 and 8. Consequently, both cases successfully regain the approximation ability. These surprising behaviors, especially the unexpected deterioration of the default choices, emphasizes the limited understanding of systematical characterizing the initialization suitable for permutation training scenarios.
@@ 4.6 Observation of the permutation-active patterns
This section aims to explore the theoretical potential of permutation training in describing network learning behavior. Based on the significant correlation between permutation and learning behavior, as evidenced by Qiu and Suda (2020) and our investigation, we hypothesize that the permutation-active components of the weights may play a crucial role in the training process. Therefore, by identifying and tracing the permutation-active part of weights, a novel description tool can be achieved, which also facilitates visualization and statistical analysis of the learning behavior.
@@
As a preliminary attempt, we illustrate the permutation behavior of the coefficients $\theta^{(2n)}$ in Fig. 2(a). The components that participated in the permutation are visually highlighted in dark green in Fig. 6(a). The behaviors plotted in Fig. 6(b)-(c) clearly show that the frequency of order relationship exchange evolves synchronously with the learning process, agreeing with the observation of Qiu and Suda (2020).
Specifically, the distribution of active components shows significant patterns, which are classified into four stages (marked by red dash lines in Fig. 6). The loss declines sharply in the initial stage, while only the components with medium value are permuted. Once loss reaches a plateau in the second stage, more components are involved in permutation, evidencing the role of permutation in propelling the training. As loss starts to decline again, the permutation frequency correspondingly diminishes. Interestingly, the slower loss decrease gives rise to a ribbon-like pattern, akin to the reported localized permutations (plotted in Fig. 14 of the appendix in Qiu and Suda (2020)), and possibly due to slow updates failing to trigger a permutation. This observation may support the existence of inherent low-dimensional structures within the permutation training dynamics, potentially linked to mathematical depiction of permutation groups, such as cycle decomposition (Cameron, 1999) and Fourier bases for permutation (Huang et al., 2009). Finally, the permutation’s saturation aligns with the stationary state of loss convergence. We believe these inspiring phenomena deserve further exploration, as they hold the potential to provide novel insights regarding the learning behavior of networks.
@@ 5.1 Generalizing to other scenarios
Although we mainly focus on basic settings, the proof idea exhibits generalizability. Extending to networks equipped with leaky-ReLU can be straightforward (refer to D for numerical evidence). Our proof also enables implementations within other designs, such as networks with deeper architectures or sparse initializations (see E for detailed discussion).
However, extending our theoretical results to the high-dimensional scenario still faces challenges. One of the primary obstacles is constructing multi-dimensional basis functions that are suitable for the permutation training scenario. A reasonable approach relies on the construction in the finite element methods (Brenner and Scott, 2008). We plan to address this problem in future work. As for the numerical evidence of high-dimensional scenarios, Qiu and Suda (2020) have examined the classification problem using VGG-based networks on the CIFAR-10 dataset, while our experiments in regression task have shown approximation behavior for two and three-dimensional inputs (see Section 4.4).
@@ 5.2 Permutation training as a theoretical tool
Our observation in Sec. 4.6 indicates the theoretical potential of permutation training, as it corresponds well with the training process and has systematical mathematical descriptions. Specifically, the patterns observed in Fig. 6 can intuitively lead to certain weight categorization strategies, potentially benefit consolidating the crucial weights for previous tasks (Maltoni and Lomonaco, 2019), or pruning to find the ideal subnetwork in the lottery ticket hypothesis (Frankle and Carbin, 2019). Additionally, the existing permutation training algorithm shares the same form as weight projection methods in continual learning (Zeng et al., 2019), as it can be viewed as applying an order-preserving projection from the free training results to the initial weight value.
@@ 5.3 The algorithmic implementation
This work is expected to facilitate the applications of permutation training. However, some practical issues still exist and deserve further investigation. As a preliminary attempt, the existing permutation training algorithm LaPerm guides the permutation by inner loops, thus incurring undesirable external computation costs. However, employing more advanced and efficient search approaches, such as the learn-to-rank formalism (Cao et al., 2007), or permutation search algorithms in the study of LMC (Jordan et al., 2023; Ainsworth et al., 2023), the benefits of permutation training will be actualized in practice. Importantly, our proof does not rely on any specific algorithmic implementations. Additionally, the incompatible initialization issue plotted in Fig. 2(b) emphasizes the need for developing more effective initializations as well as investigating the criterion of UAP-compatible initializations.